On January 1, 2015, a machine was purchased for $96,300. The machine has an estimated salvage value of $6,420 and an estimated useful life of 5 years. The machine can operate for 107,000 hours before it needs to be replaced. The company closed its books on December 31 and operates the machine as follows: 2015, 21,400 hrs; 2016, 26,750 hrs; 2017, 16,050 hrs; 2018, 32,100 hrs; and 2019, 10,700 hrs.Your answer is partially correct. Try again. Assume a fiscal year-end of September 30. Compute the annual depreciation charge
Answer #1

Machine is purchased on 1st January 2015

Cost of the machine = $ 96,300

Salvage value =6420

Useful life = 5 years

Year end is 30th September

Straight line method of depreciation-

Depreciation for each year =( Cost - salvage value) / life

= ( 96,300 - 6420 ) / 5

=$ 17,976

Year end 30th September Months of usage Calculations Depreciation amount
2015 9 months (17,976 * 9 ) / 12 13,482
2016 12 months 17,976*1 17,976
2017 12 months 17976*1 17,976
2018 12 months 17,976*1 17,976
2019 12 months 17,976*1 17,976
2020 3 months (17,976*3)/12 4,494

Sum of years digit method :

Sum of digits = n(n+1) / 2

= 5(6)/2

= 15 years

Depreciation base = cost - salvage value

= 96,300 - 6,420

= 89,880

Depreciation for 1st year (89,880)*5/15 = 29,960
Depreciation for 2nd year (89,880)*4/15 = 23,968
Depreciation for 3rd year (89,880)*3/15 = 17,976
Depreciation for 4th year (89,880)*2/15 = 11,984
Depreciation for 5th year (89,880)*1/15 = 5,992

Depreciation apportionment-

Year 2015 2016 2017 2018 2019 2020
1st year

(29,960)9/12

=22,470

(29,960)3/12

= 7,490

2nd year

(23,968 ) 9/12

= 17,976

(23,968)3/12

=5,992

3rd year

(17,976)9/12

=13,482

(17,976)3/12

=4,494

4th year

(11,984)9/12

=8,988

(11,984)3/12

=2,996

5th year

(5,992)9/12

=4,494

(5,992)3/12

=1,498

Total depreciation for the year 22,470 25,466 19,474 13,482 7,490 1,498

Double declining balance method-

Depreciation rate = (100/5)*2

= 40%

Year Book value @ the beginning (A) Calculation Depreciation amount (B) Book value at the year end (A - B = C )
2015 96,300 96,300*40%(9/12) 28,890 64,410
2016 67,410 67,410*40% 26,964 40,446
2017 40,446 40,446*40% 16,178 24,268
2018 24,268 24,268*40% 9,707 14560
2019 14,560 14560*40% 5,824 8,736
2020 8736 8736-6420 2316 0

If you are benefited from the solution then please like,if disliked then please specify the reason.